Calculate the pH for each of the following cases in the titration of 25.0 mL of 0.160 M pyridine, CsHsN(aq) with 0.160 M HBr(aq): Number (a) before addition of any HBr Number (b) after addition of 12.5 mL of HBr Number (c) after addition of 24.0 mL of HBr Number (d) after addition of 25.0 mL of HBr Number (e) after addition of 29.0 mL of HBrIL
Answer #1

Given that concentration of HBr= 0.160 M concentration of C5H5N=0.160 M Volume of C5H5N=25.0 mL Kb of C5H5N = 1.7 x 10-9 a) Bb) Addition of 12.5 mL of HBr: Number of moles of C5H5N-M*V = 0.160 M* 25.0 mL = 4.00 mmol Number of moles of HBr = M*V = 0.1c) Addition of 24.0 mL of HBr Number of moles of C5H5N = M*V = 0.160 M* 25.0 mL = 4.00 mmol Number of moles of HBr M*V 0,160d) Addition of 25.0 mL of HBr: Number of moles of C,H,N = M*V = 0.160 M* 25.0 mL = 4.00 mmol Number of moles of HBr= M*V = 0.x5.88x10 -6 a 0.08- x is small, so 0.08-x= 0.08 x-= 5.88×10 0.08 x=6.86 x 104M H1-6.86x10 M pH=-log(6.86x10*)= 3.1636795-3161